Consider the following cell reaction Fe(s)2 H ? M)-Fe2(100 M) +H>(g\1.00 atm) If the cell potential at 298 K is 0.307 volts,This is the third time I've posted this question. Can someone please help me with the top one. I slipped in a similar example. And the answer is not 2.4135
Answer #1

(1) Fe(s) + 2H+ : Fe²t + H₂ E F 24 | Fe = -0.44V; E ulty = OV E = (€° Cathocle) red. - (€ node) red. = 0 -(-0.40) EO=0.44V E=Solving for x; x²- 3. 1011 xlos Th+] = x= 5.569 xlo-3M pH = -log[ht] = -log (5.569 x103) TPH = 2.2542 (2) Ni +244 - Ni²+ +4₂